OpenStudy (anonymous):

What is the center and radius of the circle with the equation x2 + y2 – 8x + 4y – 61 = 0?

OpenStudy (anonymous):

Center: (–4, –2); Radius = 9 Center: (–4, 2); Radius = 9 Center: (4, –2); Radius = 9 Center: (4, 2); Radius = 9

OpenStudy (anonymous):

center is (-4,2) with radius 9

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

Welcome!

OpenStudy (anonymous):

What is the general form of the equation of a circle with a center at the (4, 2) and a radius of 9?

OpenStudy (anonymous):

(x – 4)2 + (y – 2)2 = 81 (x – 4)2 + (y + 2)2 = 81 (x + 4)2 + (y – 2)2 = 81 (x + 4)2 + (y + 2)2 = 81

OpenStudy (anonymous):

If you dont mind :)

OpenStudy (jennychan12):

\[(x^2-8x)+(y^2-4y) = 61\] \[(x^2-8x+16)+(y^2+4y+4) = 61+16+4\] \[(x-4)^2 +(y+2)^2 = 9^2\] general equation is \[(x+h)^2 +(y+k)^2 = r^2\] with center (h,k) and radius r.